bnmccarter:

A rectangle has sides measuring (4x + 5) units and (3x + 10) units. Part A: What is the expression that represents the area of the rectangle? Show your work. (4 points) Part B: What are the degree and classification of the expression obtained in Part A? (3 points) Part C: How does Part A demonstrate the closure property for polynomials? (3 points)

bnmccarter:

i really need some help:/

jhonyy9:

you need to know formula for area of a rectangle A = length × width

Joe348:

Part A: \[Area= Length \times Width\] \[Area= (4x+5)(3x+10)\] \[Area=12x^2+55x+50\]

jayyfrom1600:

Area of reactangle = (4x + 5)(3x + 10) ⇒ Area of recangle = 4x(3x) + 4x(10) + 5(3x) + 5(10) ⇒ Area of rectangle = 12x^2 + 40x + 15x + 50 ⇒ Area of rectangle = 12x^2 + 55x + 50 Hence, the expression which reprents the area of rectangle is 12x^2 + 55x + 50.